# Federated Variational Inference for Bayesian Mixture Models

### v1

@@ Appendix G Global Merge - Variable Selection
When finding a global clustering structure in the model with variable selection, our focus is on updating cluster allocation via merging local clusters, so we use the same global ELBO function as in the model without variable selection and only update $Z,\pi,\phi$. The nature of $\phi$ updates in the model with variable selection means that the $\epsilon^{\ast}$ parameters are equal to the prior parameters for the given covariate if a covariate is irrelevant in the model, so variable selection is still involved with the usual global merge move. Implementing this amounts to a small approximation in the true ELBO for the true Bayesian model.
We consider a variable to be ‘selected’ based on the percentage of times a variable was selected across the batches. In later simulations, we consider a variable selected if it was selected in either all batches, or all but one batch. This borrows ideas of ‘thresholds’ for variable selection across multiple VI runs which is seen in Rao & Kirk (2024).
Appendix H Approximations and Priors in MerDel
@@
In Appendix E, we detailed that in the process of performing our merge/delete moves, we ‘remove’ the clusters by removing all parameters and responsibilities associated with this cluster in the model code. This could raise concerns pertaining to transdimensional inference, as on the surface, we are deleting a dimension of our model and the variational posterior now has a different dimension.
However, we still implicitly account for the removed clusters in subsequent calculations. We call these removed clusters ‘zombie clusters’; despite not technically remaining in the model in the code, they are never truly removed but instead ruled only by priors. This is because the cluster-specific parameters $\alpha,\epsilon$ are determined by the prior only (as there is no data assigned to the cluster to update the cluster-specific parameters - see Section H.1). We can take these ‘zombie clusters’ into account in subsequent iterations of the algorithm and accurately calculate eg. ELBO values at each variational EM step by tracking the number of ‘zombie clusters’ we have.
There is an approximation made where we set responsibilities associated with ‘zombie clusters’ to 0 (by removing the responsibilities associated with the cluster). These responsibilities are not usually exactly 0. However, we describe below how the responsibilities removed are infinitesimally small for the empty cluster $k$ for all $n$ and no observations are assigned to the cluster.
H.1 Emptying a Cluster + M Update
@@
In both the merge and delete settings, when a cluster $k^{\ast}$ is removed, this is equivalent to assigning $r_{nk^{\ast}}=0$ for all observations $n$, as $r_{nk^{\ast}}$ only appears in future calculations within sums. Therefore, we can think of our $N\times K_{\mbox{init}}$ responsibility matrix as now having values of 0 in column $k^{\ast}$. After that, in both moves, we then perform a variational M step, where cluster-specific parameters $\alpha^{\ast},\epsilon^{\ast}$ are updated:
@@
As we have that $r_{nk}=0$ for all $n=1,...,N$ for a removed cluster $k$, $\alpha^{\ast}_{k}=\alpha_{0}$ and $\epsilon^{\ast}_{kjl}=\epsilon_{jl}$. These variational posteriors for cluster $k$ are solely ruled by the priors for $\pi$ and $\Phi$.
H.2 The Approximation: E Update
In MerDel, we then perform a variational E update with the full dataset (and another variational M update) to allow observations to move between clusters after the cluster-specific parameter updates. Recall that responsibilities $r_{nk}$ are defined by:
@@ H.2 The Approximation: E Update
For $x>0$, the digamma function $\psi(x)$ increases approximately in an exponential fashion.
@@
For every cluster with observations assigned to it, we generally see that the observations are assigned to it with responsibilities close to 1. Looking at Equation (53), we see that $\alpha^{\ast}_{k}$ is often a high value approximately on the scale of the number of observations assigned to the cluster - sometimes in the thousands for our dataset sizes - and the value of $\alpha_{0}$, which is chosen to be smaller than 1, has a very small impact. $\alpha^{\ast}_{k}$ must be larger for any cluster with observations compared to any ‘zombie cluster’ as $\alpha^{\ast}_{k}$ is a sum of non-negative responsibilities added to $\alpha_{0}$. $\epsilon^{\ast}$ values are similar. Therefore, from Equations (55) and (56), $\ln\rho_{nk}$ is much smaller for ‘zombie clusters’. $\rho_{nk}$ is exponentially small in comparison to other cluster, and we find that after the normalising these values, $r_{nk}$ values are extremely close to 0.
@@
It is also almost impossible that these $r_{nk}$ values will ever become larger as we continue through the EM steps; $\alpha^{\ast}$ and $\epsilon^{\ast}$ still remain extremely small for removed clusters (virtually equal to the values of the prior) in comparison to clusters with observations, and $r_{nk}$ values still remain virtually 0, and the cycle continues. Our approximation is equivalent to setting these $r_{nk}$ values to 0 permanently by removing these columns, allowing for computational efficiency and lower memory usage. We also do not perform inference for any further cluster-specific parameters associated with cluster $k$ (by removing these values too) as they are always the same as their prior values if we set $r_{nk}=0$. This saves wasted computation on these values. This approximation has virtually no effect on our model; we run two small simulations comparing the effects of implementing this approximation, and saw that true $r_{nk}$ values were often on the scale of $10^{-70}$ to $10^{-100}$ on typical MerDel example datasets (see Section H.2.1).
We then go through and perform another M step as in the previous section, before evaluating the ELBO as detailed in the next section. Going forward with the algorithm after an accepted MerDel proposal, the variational E and M steps and ELBO calculations continue as in this appendix, where $r_{nk}$ values are set to be exactly 0 forever, and we remove all values associated with removed clusters.
This is in line with what we see in the overfitted variational model without merge/delete; when clusters are ‘emptied’, they are usually never filled again with observations. When there are no observations assigned to a cluster, $\alpha^{\ast}$ and $\epsilon^{\ast}$ values are given by their prior values, and we see that $\ln\rho_{nk}$ is again much smaller for a cluster with no observations assigned to it compared to clusters with observations, and $r_{nk}$ is close to 0. Under the overfitted model, it is virtually impossible that these clusters are ever filled again, and $r_{nk}$ values will remain extremely close to 0 as illustrated above.
H.2.1 Simulation in more detail
We set up 10 simulated datasets, and compared two versions of MerDel. We performed 5 runs of the ‘fast’ version of MerDel (removing columns associated with zombie clusters) and 5 runs of the ‘slow’ version of MerDel (not removing columns associated with zombie clusters) on each simulated dataset. We used the same initialisations for the ‘fast’ and ‘slow’ versions (by setting the same random number seed). For our first simulation, each dataset had $N=2000$, $P=100$, 10 true clusters and was initialised with 15 maximum clusters, laps = 5 and $\alpha_{0}=0.01$. For the second simulation, each dataset had $N=10000$, $P=100$, 10 true clusters and was initialised with 15 maximum clusters, laps = 1 and $\alpha_{0}=0.01$.
We measured the RMSE between the final $r_{nk}$ matrices (inserting 0’s for ‘missing’ columns in the ‘fast’ MerDel version) and ARI between the final cluster labels. In both simulations, all ARIs between corresponding ‘fast’ and ‘slow’ clustering structures were 1 (perfect similarity). In the first simulation, the RMSE between $r_{nk}$ matrices ranged between $7.15\times 10^{-84}$ and $4.32\times 10^{-80}$ (with a mean of $6.63\times 10^{-81}$); in the second simulation, the RMSE ranged between $2.85\times 10^{-83}$ and $2.20\times 10^{-80}$ (with a mean of $2.78\times 10^{-81}$). When only considering $r_{nk}$ values for non-empty clusters, the RMSE was 0. The time taken for the ‘slow’ runs was indeed slower; for example, in the second simulation, the ‘fast’ runs took a mean time of 88.2 seconds to converge, while the ‘slow’ runs took a mean time of 124.2 seconds. Removing ‘zombie’ clusters saves computational time and memory usage associated with these parameters, and still gives exactly the same results.
Similar calculations show that setting $r_{nk}=0$ for removed clusters in MerDel with variable selection results in a similarly infinitesimally small approximation.
H.3 ELBO calculation
We now look at each individual term in the ELBO. We set $r_{nk}=0$ for all $n$ in removed clusters $k$ and we show that are able to accurately calculate the ELBO after completely removing all values associated with cluster $k$ from the model under this approximation.
H.3.1 Terms ruled by $r_{nk}$
$\mathbb{E}_{Z,\Phi}[\ln p(X|Z,\Phi)]$ $=\sum_{n=1}^{N}\sum_{k=1}^{K}r_{nk}(\sum_{i=1}^{D}\mathbb{E}_{\Phi}[\ln\phi_{kix_{ni}}])$ (57)
$\mathbb{E}_{Z,\pi}[\ln p(Z|\pi)]$ $=\sum_{n=1}^{N}\sum_{k=1}^{K}r_{nk}\mathbb{E}_{\pi}[\ln\pi_{k}]$ (58)
$\mathbb{E}_{Z}[\ln q(Z)]$ $=\sum_{n=1}^{N}\sum_{k=1}^{K}r_{nk}\ln r_{nk}$ (59)
In all terms (57), (58), (59) all terms in the sums are multiplied by $r_{nk}$; so all terms are equal to 0 for a removed cluster $k$. Removing terms associated with $k$ has no effect.
H.3.2 Terms ruled by $\Phi$
$\mathbb{E}_{\Phi}[\ln p(\Phi)]$ $=\sum_{k=1}^{K}\sum_{i=1}^{D}(-\ln B(\epsilon_{ki})+\sum_{l=1}^{L_{i}}(\epsilon_{kil}-1)\mathbb{E}_{\Phi}[\ln\phi_{kil}])$ (60)
$=\sum_{k=1}^{K}\sum_{i=1}^{D}\mathbb{E}_{\Phi}[\ln p(\Phi_{kj})]$ (61)
$\mathbb{E}_{\Phi}[\ln q(\Phi)]$ $=\sum_{k=1}^{K}\sum_{i=1}^{D}(-\ln B(\epsilon^{*}_{ki})+\sum_{l=1}^{L_{i}}(\epsilon^{*}_{kil}-1)\mathbb{E}_{\Phi}[\ln\phi_{kil}])$ (62)
$=\sum_{k=1}^{K}\sum_{i=1}^{D}\mathbb{E}_{\Phi}[\ln q(\Phi_{kj})]$ (63)
For the terms associated with $\Phi$, each cluster $k$ is treated independently of one another. Each $p(\Phi_{kj})$ prior and $q(\Phi_{kj})$ variational posterior is a Dirichlet distribution of dimension $L_{j}$. We can therefore look at each $(k,j)$ pair separately. For a removed cluster $k$, the $\epsilon^{\ast}_{kj}$ values associated with that cluster are equal to the prior $\epsilon_{j}$ values (Section H.1). Therefore we have, for a removed cluster $k$, $\mathbb{E}_{\Phi}[\ln p(\Phi_{kj})]-\mathbb{E}_{\Phi}[\ln q(\Phi_{kj})]=0$, as $\epsilon^{\ast}_{kjl}=\epsilon_{j}$ for all $k=1,...,K,j=1,...,P,l=1,...,L_{j}$. Ignoring cluster $k$ in both the prior and variational posterior for $\Phi$ therefore has no effect on the value of the ELBO.
H.3.3 Terms ruled by $\pi$
As in the previous section with $\Phi$, for an cluster $k$, we have that $\alpha^{\ast}_{k}=\alpha_{k}=\alpha_{0}$. Therefore, we have that the terms within the sums $\sum_{k=1}^{K}(\alpha_{k}-1)\mathbb{E}_{\pi}[\ln\pi_{k}]$ and $\sum_{k=1}^{K}(\alpha^{\ast}_{k}-1)\mathbb{E}_{\pi}[\ln\pi_{k}]$ associated with a removed cluster cancel each other out. This part of $\mathbb{E}_{\pi}[\ln p(\pi)]$ and $\mathbb{E}_{\pi}[\ln q(\pi)]$ can be considered safely dealt with.
However, $\pi$ is a $K$-length vector modelled as a $K$-length Dirichlet distribution - both under the prior and in the variational posterior. When calculating $\ln B(\alpha)$, $\ln B(\alpha^{\ast})$ and $\mathbb{E}_{\pi}[\ln\pi_{k}]$, we must take ‘zombie clusters’ into account.
$-\ln B(\alpha)=\sum_{k=1}^{K}\ln\Gamma(\alpha_{k})-\ln\Gamma(\sum_{k=1}^{K}(\alpha_{k}))$ (66)
$\ln B(\alpha^{\ast})=-\sum_{k=1}^{K}\ln\Gamma(\alpha^{\ast}_{k})+\ln\Gamma(\sum_{k=1}^{K}(\alpha^{\ast}_{k}))$ (67)
We have that $\ln\Gamma(\alpha_{k})$ and $\ln\Gamma(\alpha^{\ast}_{k})$ inside the sums cancel each other out for a removed cluster $k$, as $\alpha_{k}=\alpha^{\ast}_{k}$. Therefore we need to take into account some $\alpha,\alpha\ast$ terms in $\ln\Gamma(\sum\alpha_{k}))$, $\ln\Gamma(\sum\alpha^{\ast}_{k}))$ which do not cancel out with each other. This is easily done by simply adding $\alpha_{0}\times K_{\mbox{rem}}$ into both sums over the $\alpha$, $\alpha^{\ast}$ vectors, where $\alpha_{0}$ is the parameter in the symmetric Dirichlet prior used for $\pi$, and $K_{\mbox{rem}}$ is the number of clusters removed via merge/delete moves. This can be tracked trivially. A similar process must be used in calculating $\mathbb{E}_{\pi}[\ln\pi_{k}]$.
H.4 Approximations and Priors in FedMerDel
@@
There is no approximation when considering a ‘global merge’ in FedMerDel, as we never perform any variational E or M steps after performing merges. Our responsibilities $r_{nk}$ remain exactly 0 for ‘zombie clusters’: given that we take the zombie clusters into account when calculating the ‘global ELBO’ as above, we are okay to ‘remove’ empty clusters from the model parameters.
@@
Note that as we enter the global setting, we define a new Dirichlet prior for $\pi$ over our full dataset with the total number of local clusters initialised: $B\times K_{b}$, if $K_{b}$ is the initialised number of clusters in Batch $b$. This is justified as each cluster has a prior where:
@@
with K representing the number of maximum clusters desired in the model. Therefore, when we enter the global setting with $B$ batches, the maximum clusters in the model should be $B\times K_{b}$, the maximum total clusters when we consider all batches. A fractional prior for each batch would not be appropriate; this would be imposing a prior saying that there is a maximum number of $K_{b}$ clusters across the full dataset, despite the fact we actually search for this number of clusters in each partition. We do not want to constrain the model to force all clusters to merge across batches; especially in the federated learning setting, there may be clusters appearing in certain batches of data but not other batches. This is consistent with similar literature; the same Dirichlet priors for mixing probabilities are used in ‘local clusters’ and ‘global clusters’ in (Zuanetti et al., 2018), a similar MCMC model to FedMerDel.
Appendix I Simulation Set-Up - Additional Information
I.1 Data Generating Mechanism
@@
I.2 Hyperparameter Choice
In all simulations (real-world and simulated data), we set $\alpha_{0}=0.01$, $\epsilon_{j}=1/L_{j}$ and $a=2$ (as in Rao & Kirk (2024)) in the priors in Section 2 of the main paper. We did not tune these in this manuscript as all examples are purely illustrative. All algorithms were run with a maximum of 1000 iterations (an iteration being one variational EM step) - although this limit was not reached for any simulation.
Convergence tolerance was set between 0.000005 and 0.00000005 in all simulations. When comparing wall-clock time for convergence, all simulations in a given study had the exact same convergence tolerance.
In all simulations involving FedMerDel, we used the parameters ‘laps = 5’ for MerDel runs, except when explicitly specified.
I.3 Frequency of Merge/Delete Moves
Details of the scenarios for this simulation study are given in Table 2 Simulations 1.1, 1.2, 1.3 were run without variable selection; Simulations 1.4, 1.5 were run with variable selection. $K_{\mbox{init}}$ refers to the initialised value of $K$ in the model. In each scenario, we generated 20 independent datasets and compared 10 different initialisations for each dataset.
@@ I.3 Frequency of Merge/Delete Moves
ID Relevant $N$ $P$ $K_{init}$ $K_{true}$ $N$ per
@@
I.4 Global Merge Simulations
We aimed to evaluate the clustering performance of FedMerDel when splitting a simulated dataset into equally sized batches, and compared this to a paralellised version of MerDel (on the full dataset) as well as the usual MerDel algorithm on the full dataset (usual variational EM with merge and delete moves). We set ‘laps = 5’, $P=100$ covariates, 12 true clusters and 20 initialised clusters in all cases. Correlation was used for all merge criteria. These parameters were all kept the same to enable comparison of run-times between different values of $N$. Details of the values of $N$ tested and number of batches/cores used are given in Table 3.
@@ I.4 Global Merge Simulations
Table 3: Table giving parameters for data generation for the second simulation study, ‘Global Merge Simulations’, where other parameters are set as in Section I.4.
@@
$\mathbf{K}$=6 simulations
We looked at some simulations with a lower number of ‘true clusters’ (6). All datasets had $P=100$ covariates, and we split these into 20 batches or parallelised over 20 cores. We initialised with 10 clusters in all MerDel cases. We compare FedMerDel to parallelised MerDel. We looked at $N=100,000$ and $N=500,000$.
@@
We additionally compared the ARI, number of clusters and the time taken for clustering models using parallelised MerDel and FedMerDel when we varied $P$. In this simulation, we simulated 5 synthetic datasets and ran MerDel and FedMerDel with 5 ‘shuffles’/initialisations each time. Datasets were all of size $N=100,000$ with 10 equally sized clusters, and we initialised with 20 clusters in all cases. We split into 5 equally sized batches for FedMerDel, and parallelised over 5 cores for parallelised MerDel. Results can be found in Appendix J.2.2.
@@
For this simulation with variable selection (on binary data), we employed the ‘greedy search’ for the global merge, and used correlation to assess similarity between clusters. The two scenarios had simulated data split into 5 equally sized batches with $N=$ 20,000, 50,000, with $K=$ 15 and 10 equally sized true clusters respectively. We set $P=100$ covariates in both cases, where 80% were relevant to the clustering structure in the first simulation and 75% were relevant in the second simulation. We generated 5 simulated datasets and had 5 different ‘shuffles’ of the data, as well as running MerDel on the full data 5 times (unparallelised and parallelised), as with other ‘global merge’ simulations in this manuscript.
As well as looking at ARI, number of resulting clusters and wall-clock time as with other simulations, we looked at the number of selected variables. As described in Section G, for FedMerDel, we consider a variable selected if it was selected in either all batches, or all but one batch (where a variable is selected if $c_{j}>0.5$, the expectation of the variable selection latent variable (Appendix D)).
I.5 Number of Batches

### v2

@@ D.3 Global Merge - Variable Selection
When finding a global clustering structure in the model with variable selection, our focus is on updating cluster allocation via merging local clusters, so we use the same global ELBO function as in the model without variable selection and only update $Z,\pi,\phi$. Implementing this amounts to a small approximation in the true ELBO for the true Bayesian model. We consider a variable to be ‘selected’ based on the percentage of times a variable was selected across the batches. In later simulations, we consider a variable selected if it was selected in either all batches, or all but one batch. This borrows ideas of ‘thresholds’ for variable selection across multiple VI runs, seen in Rao2024.
Appendix E Approximations and Priors in MerDel
@@
In Appendix C, we detailed that in the process of performing our merge/delete moves, we ‘remove’ the clusters by eliminating their parameters and responsibilities from the model code. This could raise concerns pertaining to transdimensional inference. However, we implicitly account for the removed clusters in subsequent calculations. We call these removed clusters ‘zombie clusters’.
This is because the cluster-specific parameters $\alpha,\epsilon$ are determined by the prior only (as there is no data assigned to the cluster to update cluster-specific parameters). We can accurately calculate ELBO values and track these zombie clusters throughout the algorithm. The key approximation is setting responsibilities for zombie clusters to exactly 0. These values are not usually precisely 0, but are infinitesimally small.
E.1 Emptying a Cluster + M Update
@@
When a cluster $k^{\ast}$ is removed (through merge or delete), this is equivalent to assigning $r_{nk^{\ast}}=0$ for all observations $n$, effectively zeroing column $k^{\ast}$ in the $N\times K_{\mbox{init}}$ responsibility matrix. The subsequent variational M step updates $\alpha^{\ast},\epsilon^{\ast}$:
@@
As we have that $r_{nk}=0$ for all n=1,,Nn=1,...,N for a removed cluster $k$, $\alpha^{\ast}_{k}=\alpha_{0}$ and $\epsilon^{\ast}_{kjl}=\epsilon_{jl}$. These variational posteriors for cluster $k$ are solely ruled by the priors for $\pi$ and $\Phi$.
E.2 The Approximation: E Update
In MerDel, we then perform a variational E update with the full dataset to allow observations to move between clusters after cluster-specific parameter updates. Recall that responsibilities $r_{nk}$ are defined by:
@@ E.2 The Approximation: E Update
Since, for $x>0$, the digamma function $\psi(x)$ increases approximately exponentially, observations are assigned to clusters with responsibilities close to 1. These clusters therefore have $\alpha^{\ast}_{k}$ values on the scale of the number of observations assigned - sometimes in the thousands - which are much larger than $\alpha_{0}<1$. Therefore $\ln\rho_{nk}$ is much smaller for ‘zombie clusters’, resulting in $r_{nk}$ values extremely close to 0.
@@
These near-zero $r_{nk}$ values persist as we continue through the EM steps; $\alpha^{\ast}$ and $\epsilon^{\ast}$ still remain extremely small for removed clusters (virtually equal to prior values) in comparison to clusters with observations, and $r_{nk}$ values still remain virtually 0. Our approximation sets $r_{nk}=0$ permanently by removing these columns, improving computational efficiency and memory usage. Since cluster-specific parameters for removed clusters equal their priors when $r_{nk}=0$, we also remove these.
This approximation has virtually no effect on our model; we run two small simulations comparing the effects of implementing this approximation, and saw that true $r_{nk}$ values were often on the scale of $10^{-70}$ to $10^{-100}$ on typical MerDel datasets. After an accepted MerDel proposal, we continue with standard variational EM steps where removed clusters remain permanently eliminated.
This behaviour mirrors the overfitted variational model without merge/delete, where empty clusters naturally maintain near-zero responsibilities and virtually never regain observations. ELBO calculation remains straightforward; terms from zombie clusters either cancel each other out or are trivially calculated.
E.3 Approximations and Priors in FedMerDel
@@
There is no approximation when considering a ‘global merge’ in FedMerDel, as we never perform any variational E or M steps after performing merges. Responsibilities $r_{nk}$ remain exactly 0 for ‘zombie clusters’.
@@
In the global setting, we define a new Dirichlet prior for $\pi$ over our full dataset with the total number of local clusters initialised: $B\times K_{b}$, where $K_{b}$ is the initialised number of clusters in Batch $b$. This is justified as each cluster has a prior where:
@@
with K representing the maximum desired clusters per batch. The maximum clusters globally should be $B\times K_{b}$, the maximum total clusters when we consider all batches. A fractional prior for each batch would inappropriately constrain the model to $K_{b}$ clusters across the entire dataset, despite searching $K_{b}$ clusters per partition. We do not want to constrain the model to force all clusters to merge across batches; in FL, there are often clusters only appearing in certain batches. This is consistent with literature; the same Dirichlet priors for mixing probabilities are used in ‘local clusters’ and ‘global clusters’ in (Zuanetti2018), a similar MCMC model.
Appendix F Implementation Details
The model builds upon an existing R package, VICatMix (Rao2024), which applies VI for binary and categorical Bayesian finite mixture models (with no merge/delete moves). Rcpp and RcppArmadillo are used to accelerate computation with C++. MerDel runs on a single-core processor, where, for example, inference for a mixture model fitted to a binary dataset of size $N=2000$, $P=100$ generally converges in less than a minute.
FedMerDel allows MerDel to be run independently on a different core per batch of data; reported ‘wall-clock times’ take this parallelisation into account where clearly stated.
Appendix G Simulation Set-Up - Additional Information
G.1 Data Generating Mechanism
@@
G.2 Hyperparameter Choice
In all simulations (real-world and simulated data), we set $\alpha_{0}=0.01$, $\epsilon_{j}=1/L_{j}$ and $a=2$ (as in Rao2024) in the priors in Section 2 of the main paper. We did not tune these in this manuscript as all examples are purely illustrative. All algorithms were run with a maximum of 1000 iterations (an iteration being one variational EM step) - although this limit was not reached for any simulation.
Convergence tolerance was set between 0.000005 and 0.00000005 in all simulations. When comparing wall-clock time for convergence, all simulations in a given study had the same convergence tolerance. In all simulations involving FedMerDel, we used the parameters ‘laps = 5’ for MerDel runs, except when specified.
G.3 Frequency of Merge/Delete Moves
Details of the scenarios for this simulation study are given in Table 3. Simulations 1.1, 1.2, 1.3 were run without variable selection; Simulations 1.4, 1.5 were run with variable selection. $K_{\mbox{init}}$ refers to the initialised value of $K$ in the model. In each scenario, we generated 20 independent datasets and compared 10 different initialisations for each dataset. We used correlation for merge criteria.
@@ G.3 Frequency of Merge/Delete Moves
ID Relevant NN $P$ $K_{init}$ $K_{true}$ NN per
@@
G.4 Global Merge Simulations
We aimed to evaluate the clustering performance of FedMerDel when splitting a simulated dataset into equally sized batches, and compared this to a parallelised version of MerDel (on the full dataset) as well as the usual MerDel algorithm on the full dataset (variational EM with merge and delete moves). We set ‘laps = 5’, $P=100$ covariates, 12 true clusters and 20 initialised clusters in all cases. Correlation was used for all merge criteria. Parameters were kept the same to enable comparison of run-times between different values of $N$. Details of the values of $N$ tested and number of batches/cores used are given in Table 4.
@@ G.4 Global Merge Simulations
Table 4: Table giving parameters for data generation for the second simulation study, ‘Global Merge Simulations’, where other parameters are set as in Section G.4.
@@
Heterogeneous Data Simulations
As the above simulations have observations assigned purely randomly to batches, we also look at situations with heterogeneous (non i.i.d) data in batches, as is common in FL. We consider scenarios where a) 1 of 12 clusters is only seen in one batch and no others with $N=50,000$, b) 10 clusters are equally split between 5 batches with $N=50,000$ and c) 10 clusters are equally split between 5 batches, with 2 more clusters which are evenly split across the 5 batches with $N=20,000$. In all cases, we set $P=100$. We compared this to parallelised VI on the full data.
Comparisons to Other Unsupervised FL Methods
We compared to other unsupervised FL methods for simulated data, including FedEM for a categorical mixture model (NEURIPS2021_f740c8d9), and an adaptation of k-FED (pmlr-v139-dennis21a) that uses k-modes instead of k-means as appropriate for categorical data. We also compared an approximate version of SNOB (Zuanetti2018), with local clusters found via Gibbs sampling (with R package PReMiuM (Liverani2015)), then merged using hierarchical clustering. We looked at randomly distributed simulated data with $N=20,000$ and $N=50,0000$, and also looked at heterogeneous data scenario c) in the above paragraph.
Comparisons to Other Unsupervised Centralised Learning Methods
To provide additional context for MerDel and FedMerDel’s clustering performance, we compared MerDel and FedMerDel to centralized clustering methods suitable for large binary datasets on simulated datasets of size $N=20,000$ and $N=50,000$. Methods we looked at were k-modes (Chaturvedi2001) (recall this is also used for initialisation of MerDel), hierarchical clustering with ‘hclust’ in R (Murtagh2011), and DBSCAN (dbscan) - in all cases, we fed the algorithm the true number of clusters, 12. We also evaluated latent class analysis (LCA), which previously performed well on similar binary EHR data (Nichols2022-hq) using the poLCA R package (Linzer2011).
@@
We additionally compared the ARI, number of clusters and the time taken for clustering models using parallelised MerDel and FedMerDel when we varied $P$. In this simulation, we simulated 5 synthetic datasets and ran MerDel and FedMerDel with 5 ‘shuffles’/initialisations each time. Datasets were all of size $N=100,000$ with 10 equally sized clusters, and we initialised with 20 clusters in all cases. We split into 5 equally sized batches for FedMerDel, and parallelised over 5 cores for parallelised MerDel.
Less Separable Data
To address the challenge where clusters are less easily separated, we also ran simulations where the probability of a ‘1’ was generated by a Beta(2, 2) distribution. Further, we simulated data with this generating distribution and where two clusters - clusters 11 and 12 - had 75% of their covariates with their cluster-specific probabilities being $\pm$ 0.1 of each other. In both of these simulations, we set $N=50,000$, $P=100$ and 12 true clusters, initialised with 15 clusters. 5 independent datasets were generated, and FedMerDel was run 3 times on each.
@@
For this simulation with variable selection (on binary data), we employed the ‘greedy search’ for the global merge, and used correlation to assess similarity between clusters. The two scenarios had simulated data split into 5 equally sized batches with $N=$ 20,000, 50,000, with $K=$ 15 and 10 equally sized true clusters respectively. We set $P=100$ covariates in both cases, where 80% were relevant to the clustering structure in the first simulation and 75% were relevant in the second simulation. We generated 5 simulated datasets and had 5 different ‘shuffles’ of the data, as well as running MerDel on the full data 5 times (unparallelised and parallelised), as with other ‘global merge’ simulations.
As well as looking at ARI, number of resulting clusters and wall-clock time as with other simulations, we looked at the number of selected variables. As described in Section D.3, for FedMerDel, we consider a variable selected if it was selected in either all batches, or all but one batch (where a variable is selected if $c_{j}>0.5$, the expectation of the variable selection latent variable (Appendix B)).
G.5 Number of Batches